Find an equation of the line containing the given pair of points.
(1/6, - 1/3)and (5/6,1)
What is the equation of the line y?
(y=mx+b)

Answer by checkley77(12844)

FIRST WE FIND THE SLOPE(m)
SLOPE=(Y2-Y1)/(X2-X1)=(1+1/3)/(5/6-1/6)=(4/3)/(4/6)= (4/3)/(2/3)=4/3*3/2=12/6=2
-1/3=2*1/6+b
-1/3=1/3+b
b=-1/3-1/3
b=-2/3
THUS THIS LINE EQUATION IS:
Y=2X-2/3